OpenStudy (anonymous):

The midpoint of line UV is (5,-11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (morganchavez):

wouldnt it be 7,27

OpenStudy (morganchavez):

do you have multiple choice because that would help me more i do my math backwards lol dont ask i jus learned that way

OpenStudy (zpupster):

finding midpoint is p1+p2/2 (x1 +x2)/2 = midpoint x and (y1 +y2)/2 = midpoint y we know midpoint is 5 for x and endpoint is 3 so (3+x)/2=5 multiply both sides by 2 3+x=10 2's canceled 3-3+ x = 10 -3 x=7 do the same to figure out your y